Solve the equation
:3s -3: = :s-1:

Answer by stanbon(75887)

:3s -3: = :s-1:
|3s - 3| = |s -1|\
3|s-1| = |s-1|
For any value of s,
3(s-1) = (s-1)
3 = 1
-------------------
This is true for no values of "s".
-------------------
Cheers,
Stan H.